# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v3

@@ 4. Convergence of NGD for Two-Layer PINNs
Assumption 4.3 holds for various commonly used activation functions, including logistic function $\sigma(z)=1/(1+e^{-z})$ (with $\phi(z)=e^{-z}$), softplus function $\sigma(z)=\log(1+e^{z})$ (with $\phi(z)=e^{-z}$), hyperbolic tangent function $\sigma(z)=(e^{z}-e^{-z})/(e^{z}+e^{-z})$ (with $\phi(z)=e^{-2z}$), swish function $\sigma(z)=z/(1+e^{-z})$ (with $\phi(z)=ze^{-z}$) and others. Compared to the $\text{ReLU}^{3}$ activation function, these smooth activation functions are more popular for PINNs because solving PDEs typically requires high-order derivatives.
@@ 10. Proof of Section 4
For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary C.1, respectively.
@@
Thanks to Corollary C.1, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary C.1 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary C.1 leads to Theorem 4.7.
@@
In the proof, we assume that Condition 3 holds for $t=0,\cdots,k$. Then from Corollary C.1, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$. Here, we primarily explain the process from Condition 3 to Corollary C.1, while other content is placed in the following full proof of Theorem 4.7.
@@
where the last inequality is due to Corollary C.1.
@@
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary C.1, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ implies that

### v4

@@ 4. Convergence of NGD for Two-Layer PINNs
Assumption 4.3 holds for various commonly used activation functions, including logistic function $\sigma(z)=1/(1+e^{-z})$ (with $\phi(z)=e^{-z}$), softplus function $\sigma(z)=\log(1+e^{z})$ (with $\phi(z)=e^{-z}$), hyperbolic tangent function $\sigma(z)=(e^{z}-e^{-z})/(e^{z}+e^{-z})$ (with $\phi(z)=e^{-2z}$), swish function $\sigma(z)=z/(1+e^{-z})$ (with $\phi(z)=ze^{-z}$) and others. Although $\sin(z)$ and $\cos(z)$ do not satisfy the condition (16), it can be seen from Remark 10.2 in the appendix that the Gram matrices induced by them are also strictly positive definite. Compared to the $\text{ReLU}^{3}$ activation function, these smooth activation functions are more popular for PINNs because solving PDEs typically requires high-order derivatives.
@@ 10. Proof of Section 4
Remark 10.1.
The key point in the proof lies in the fact that the order of the PDE in the interior is higher than that of the initial and boundary conditions, allowing for a natural extension to broader classes of PDEs. For general PDEs, we may focus solely on the interior and boundary, assuming the interior is of second order and the boundary is of first order. Suppose the second-order interior term is taken at $x_{0}$, i.e., it has the form $\frac{\partial^{2}u}{\partial x_{0}^{2}}$, and the first-order boundary term is also taken at $x_{0}$. Since we can translate the coordinates, without loss of generality, we can assume that all $x_{0}$-components are positive.
For the interior, taking the $k$-th derivative of $w_{0}^{2}\sigma^{(2)}(\bm{w}^{T}\bm{x})$ yields that
$w_{0}^{2}\sigma^{(k+2)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+2kw_{0}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{1}^{(i)}\otimes\bm{x}^{\otimes(k-1)}$
$\quad+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{0}^{(i)}\otimes\bm{e}_{0}^{(j)}\otimes\bm{x}^{\otimes(k-2)}.$
For the boundary, taking the $k$-th derivative of $w_{0}\sigma^{(1)}(\bm{w}^{T}\bm{x})$ yields that
$\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)}.$
As before, we set $\bm{w}=(w_{0},\bm{0})$. Then, the equation (68) becomes
$w_{0}^{2}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k+1)}(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]+w_{0}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}^{0}\right]+\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}^{1}\right]$
$\quad+w_{0}\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{y}_{j}^{\otimes(k)}\right]+\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{z}_{j}^{2}\right]=0,$
where $\bm{z}_{p}^{0},\bm{z}_{p}^{1},\bm{z}_{j}^{2}$ are tensors of similar form $\bm{z}_{p}$ in (69), whose explicit definitions are omitted for simplicity. Dividing both sides by $w_{0}\phi(w_{0}x_{p}^{0})$ reduces it to the form considered earlier. We can therefore conclude that the Gram matrix is strictly positive definite. Indeed, since the orders of the interior and boundary terms in the partial differential equation differ, we can relax the conditions in Lemma 4.4 to simply requiring that no two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ are parallel and no two samples in $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel. In brief, we can set $k=n_{1}-1$ and $k=n_{2}-1$ in equation (67), and then use the method described above.
Remark 10.2.
For the activation functions $\sin(x)$ and $\cos(x)$, in equation (68), we may assume that $\sigma^{(k+1)}(x)=\sin(x),\sigma^{(k)}(x)=-\cos(x)$. Dividing both sides of equation (68) by $w_{0}$ and letting $w_{0}\to+\infty$, we can obtain
$\lim\limits_{w_{0}\to+\infty}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sin(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]=0.$
We express the general form of the components of the tensor above as
$\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}\sin(w_{0}x_{p}^{0})=\sum\limits_{i=1}^{n}a_{i}\sin(w_{0}b_{i}),$
where the $b_{i}>0$ are distinct and $c_{p}$ denotes the components of the tensor $\bm{x}_{p}^{\otimes(k)}$. For simplicity, we denote $w_{0}$ as $x$. To prove that $\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0$, we need to show that any component of this tensor is zero, i.e., its general form satisfies $\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}=0$. This is equivalent to proving $\sum\limits_{i=1}^{n}a_{i}=0$.
Let $f(x)=\sum\limits_{i=1}^{n}a_{i}\sin(b_{i}x)$, note that dividing both sides of equation (68) by $w_{0}$ yields that $f(x)=\mathcal{O}(1/x)$. Thus, we can consider the average energy of $f^{2}(x)$ over the interval $[T,T+L]$, i.e.,
[…]
Expanding this, we obtain
[…]
$=\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}-\frac{1}{L}\sum\limits_{i=1}^{n}\frac{\sin(2b_{i}(T+L))-\sin(2b_{i}T)}{4b_{i}}$
$\quad+\frac{1}{L}\sum\limits_{i\neq j}a_{i}a_{j}\frac{\sin((b_{i}-b_{j})(T+L))-\sin((b_{i}-b_{j})T)}{2(b_{i}-b_{j})}$
$\quad-\frac{1}{L}\sum\limits_{i\neq j}a_{i}a_{j}\frac{\sin((b_{i}+b_{j})(T+L))-\sin((b_{i}+b_{j})T)}{2(b_{i}+b_{j})}.$
Taking the limits $L\to+\infty$ and $T\to+\infty$ in the above equation, the right-hand side tends to $\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}$. Regarding the left-hand side, recall that $f(x)=\mathcal{O}(1/x)$, thus for any $\epsilon>0$, there exists $T_{0}$ such that for all $x>T_{0}$, $|f(x)|<\epsilon$. Therefore, for $T>T_{0}$ and any $L$, we have
[…]
By the arbitrariness of $\epsilon$, we can deduce that
[…]
Hence, we can deduce that $\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}=0$, which implies that $\sum\limits_{i=1}^{n}a_{i}=0$, i.e., $\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}=0$. Finally, we obtain
$\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0.$
Applying the same approach as before, we conclude that $\alpha_{p}=0$ for all $p\in[n_{1}]$and $\beta_{j}=0$ for all $j\in[n_{2}]$.
@@
For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary 10.3, respectively.
@@
Thanks to Corollary 10.3, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary 10.3 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 10.3 leads to Theorem 4.7.
@@
In the proof, we assume that Condition 3 holds for $t=0,\cdots,k$. Then from Corollary 10.3, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$. Here, we primarily explain the process from Condition 3 to Corollary 10.3, while other content is placed in the following full proof of Theorem 4.7.
@@
where the last inequality is due to Corollary 10.3.
@@
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary 10.3, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ implies that